OpenStudy (anonymous):

someone please help me with this acceleration problem...A truck travels at 40 m/s and slows down to 10 m/s in 5 seconds. What is the acceleration of the truck? Show formula

OpenStudy (anonymous):

v = u + at Here u = 40 m/s v = 10 m/s t= 5 s Let acceleration = A substitute 10 m/s = 40 m/s + A * 5 s (10 - 40) m/s = A * 5 s - 30 m/s = A * 5 s (-30/5) m/s^2 = A So A = -6 m/s^2 The negative sign shows that it is a negative acceleration ( as truck has solowed down).

OpenStudy (anonymous):

acceleration = (final velocity - inicial velocity) / time so: a=10-40/5 a=-6m/s